OpenStudy (annie96):

1.)The graph of a line has a slope of -3 and passes through the point (-6,-4). What is the equation of the line in point-slope form? y+4=-3(x+6) y-4=-3(x-4) y-6=-3(x-4) y+6=-3(x+4) 2.)What are the x- and y-intercepts of the line? 2x-3y=-12 x-intercept 6, y-intercept -4 x-intercept-6, y-intercept 4 x-intercept-4, y-intercept 6 x-intercept 4, y-intercept -6

OpenStudy (anonymous):

@Annie96 ...The standard form of a straight line in point-slope form is: \[y-b=m(x-a)\] for a point (a,b) and slope m...Can you plugin you values above and show me your work?

OpenStudy (annie96):

2(6)-3(-4)=12

OpenStudy (anonymous):

Nope...for your first question, point=(a,b)=(-6,-4) and slope=m=-3 so plugin these values in the standard equation of straight line: \[y-b=m(x-a)\]

OpenStudy (annie96):

-4+4=-3(-6+6)

OpenStudy (anonymous):

no you're putting (x,y)=(-6,-4) as well...they're general variables...(x,y) could be any point on the line... it should be : y+4=-3(x+6)

OpenStudy (annie96):

ok.

OpenStudy (anonymous):

If you want to know more about how STANDARD FORM of equation for straight line is derived: http://www.mathsisfun.com/algebra/line-equation-point-slope.html or you could post another question

OpenStudy (anonymous):

About your second question: Hint: i) X intercept is the value of x at which the line cuts the x-axis i.e y=0 for X-intercept ii) Y intercept is the value of y at which the line cuts the y-axis i.e x=0 for Y-intercept so can you guess your x and y intercepts

OpenStudy (annie96):

x-intercept-6, y-intercept 4

OpenStudy (anonymous):

That's right

OpenStudy (annie96):

Thank you so much for you're help.

OpenStudy (anonymous):

oh no problem